Suppose during volleyball practice, you lost 2.0 lbs of water due to sweating. If all of this water evaporated, how much energy did the water absorb from your body? Express your answer in kJ.  2.2 lbs = 1.0 kg


1
Expert's answer
2020-11-02T08:52:37-0500

Given

2.2 lbs = 1.0 Kg

2.0 lbs = "\frac{1.0}{2.2}" x 2.0 = 0.91 kg = 910 g


Q = n x ΔH

Q = heat gained

n = no. of moles = "\frac{Given.mass}{molecular.mass}"


n = no. of moles = "\frac{910g}{180g/mol}" = 50.55 moles


ΔH = enthalpy of vaporization of water

ΔH = 40.65 KJ/moles


Q = n x ΔH

Q = 50.55 moles x 40.65 KJ/moles

Q = 2054.86 KJ